can someone explain how to solve for kp assuming this datat

\lambda F(\lambda) \sigma(\lambda ) \phi(\lambda )
295
301
305
311
315
319
332.5
337.5
342.5
347.5
352.5
357.5
362.5
372.5
377.5
382.5
387.5
392.5
397.5
402.5
407.5
412.5
0
0.01
0.09
0.61
1.13
1.7
10.22
10.14
11.12
11.33
13.09
11.84
13.23
1.48
1.69
1.44
1.59
1.6
1.97
2.74
2.68
2.8
11.54
13.44
15.89
18.67
24.53
24.77
34.25
37.98
37.98
40.65
43.13
47.17
48.33
53.15
55.08
56.44
59.27
58.45
60.21
57.81
59.99
56.51
0.998
0.997
0.996
0.995
0.994
0.993
0.991
0.99
0.98
0.988
0.987
0.946
0.984
0.981
0.979
0.975
0.996
0.957
0.87
0.56
0.225
0.094

kp= \sum   \phi (\lambda )\sigma (\lambda )F(\lambda )

  \lambda
Answer #1

Just multiply F(λ)Φ(λ)σ(λ) at each λ and add all those values,
At λ = 295 , KP = 0.998*11.54*0 = 0
At λ= 301 , KP = 0.997*13.44*0.997 = 0.133996
At λ= 305 , KP = 0.996*15.89* 0.09 = 13.346081
At λ = 311, KP = 0.995*18.67*0.61 = 11.3317565
At λ = 315, KP = 0.994*24.53*1.13 = 27.5525866
At λ = 319,  KP = 0.993*24.77*1.7 = 41.814237
At λ = 332.5 , KP = 0.991*34.25*10.22 = 346.884685
At λ = 337.5 , KP = 0.99*37.98*11.12 = 418.114224
.
.
.
At λ= 412.5 , KP = 0.094*56.51*2.8 = 14.873432
Just add all these values, to get KP
KP = 4067.92366